OpenStudy (anonymous):

Amy has a box containing 6 white, 4 red, and 8 black marbles. She picks a marble randomly. It is red. The second time, she picks a white marble. In the third attempt, Amy picks a white marble again. Amy has not replaced the marbles she picked. The probability that the fourth marble Amy picks is black is ........?. If Amy picks a black marble in the fourth attempt, the probability that in the fifth attempt she will pick a red or a black marble is ........?.

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

this one seems to just be a longer drawn out version of the last how many balls to start with?

OpenStudy (anonymous):

15? well that's without replacement

OpenStudy (amistre64):

18-3 = 15 good, and there are still 8 black eyes left to pick from right?

OpenStudy (anonymous):

eyes? haha but yea so P(black)=8/15 right?

OpenStudy (amistre64):

marbles have cateyes :) yes, the probability that the 4th pick is a black marble given the prior results is: 8 out of 15

OpenStudy (anonymous):

what?! no way... haha wtf xD you serious???.... and alright i jus need the fraction, by the way..

OpenStudy (amistre64):

now there are 14 marbles left ... how many left are red or black?

OpenStudy (anonymous):

3 red, 7 black, 4 white

OpenStudy (amistre64):

3+7= 10 that make a favored outcome whats the probability of picking 10 out of 14 ?

OpenStudy (anonymous):

so reduced its 5/7 ... right?

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

sweet alright 3 more... one moment

OpenStudy (anonymous):

thanks for the help by the way, i really appreciate it

OpenStudy (amistre64):

yep, as long as youre participating in the solution process, im fine with it